OpenStudy (anonymous):

you have a savings account that offer an annual simple interest rate of 5%. You put $350 into the account. Draw a graph showing the total simple interest you earn after 1 ,2 ,3 and 4 years

OpenStudy (anonymous):

add $17.50 a year i think

OpenStudy (anonymous):

can you draw a graph please

OpenStudy (anonymous):

I need this I was stuck here for HOURS

OpenStudy (anonymous):

@jdoe0001 can you help please

OpenStudy (jdoe0001):

\(\large {i=prt \\ \quad \\ p=principal\to 350\\ r=rate\to 5\%\to \frac{5}{100}\to 0.05\\ t=years\to \{x:1,2,3,4\} \\ \quad \\ i=350\cdot 0.05\cdot t\implies y=350\cdot 0.05\cdot {\color{red}{ x}} \\ \quad \\ \begin{array}{llll} x\qquad &y \\\hline\\ 1&y=350\cdot 0.05\cdot {\color{red}{ 1}}\\ 2&y=350\cdot 0.05\cdot {\color{red}{ 2}}\\ 3&y=350\cdot 0.05\cdot {\color{red}{ 3}}\\ 4&y=350\cdot 0.05\cdot {\color{red}{ 4}} \end{array} }\) get the "x" and "y" values and graph away

OpenStudy (anonymous):

the first one was right